# Apparent delay of the Kibble-Zurek mechanism in quenched open systems

### v1

@@ V Summary and Discussion
In this work, we extend the Kibble-Zurek mechanism to open systems under a finite quench and report an intermediate regime separating the breakdown and validity of the KZM at fast and slow quench timescales, respectively. This novel regime manifests as a continuation of the transition time’s KZM power-law scaling at $\tau_{q}$ where the system appears to relax after the quench has terminated. As we have shown using a coupled oscillator system, this phenomenon results from the system’s relaxation towards the global minimum of its potential due to dissipation. This mechanism effectively hides the system’s crossover to the adiabatic regime, only to be revealed once the system reaches the arbitrary threshold of the order parameter.
Using the open DM and the open DLM, we have also demonstrated that the delayed KZM is a generic feature of open systems under finite quenches that can be mapped onto a coupled oscillator system. Furthermore, we have shown that the signatures of the delayed KZM, specifically the size of the quench interval where the delayed KZM regime is observed, become more prominent for small values of $\varepsilon_{f}$, highlighting the dissipative and near-critical nature of this phenomenon. We have discussed the implications of the delayed KZM in the context of the threshold-based criterion typically used in experiments to measure the transition time and proposed an alternative method to measure $\hat{t}$. Our proposed method only relies on the spatial-temporal information of an appropriate order parameter, such as the defect number and phase information of the system’s bosonic modes, thus providing a more sensitive tool for detecting phase transitions.
Our results extend the notion of the KZM to dissipative systems with finite quench protocols beyond the limits of slow and rapid quenches. It also provides a framework on how the manifestation of the KZM can be altered in experimental protocols, wherein limitations in measuring the true AI crossover become more relevant. Since our results are all in the mean-field level, a natural extension of our work is to verify whether the delayed KZM would survive in the presence of quantum fluctuations. It would also be interesting to test the signatures of the delayed KZM in the quantum regime of the open DM and the open DLM, and further explore their universality classes beyond the mean-field level. These extensions can be readily done in multiple platforms, including, but not limited to, cavity-QED set-ups [55, 56, 57, 27, 40], nitrogen-vacancy centre ensembles [43, 58, 59, 60], cavity-magnon systems [45, 46], and photonic crystals [61].
Acknowledgement
@@ Appendix A Mean-field Equations of the Considered Systems
To obtain the mean-field equations of the systems considered in the main text, we consider the master equation for the expectation value of an arbitrary operator, $\hat{O}$
@@
where $\hat{H}$ is the system’s Hamiltonian, and $\mathcal{D}\hat{O}=\sum_{\ell}\kappa_{\ell}\left(2\hat{L}_{\ell}^{\dagger}\hat{O}\hat{L}_{\ell}-\left\{\hat{L}^{\dagger}_{\ell}\hat{L}_{\ell},\hat{O}\right\}\right)$ is the dissipator, with $\hat{L}_{\ell}$ being the jump operators. We will also let $A=\left<\hat{A}\right>$ for notation convenience. Using this master equations, the mean-field equation of the open COS is,
@@
As for the open Dicke model, its mean-field equations are,
@@ Appendix B KZM exponents of the open Dicke Lattice model
where $D$ and $d$ are the dimensions of the system and the topological defects, respectively. To demonstrate that the open DLM satisfies the predicted KZM scaling for $N_{d}$, we present in Fig. 9 the number of phase slips present in the system for a given $\tau_{q}$ and $\varepsilon_{f}$. We can observe that for large $\tau_{q}$, $N_{d}$ follows a power-law scaling behaviour with $\tau_{q}$, emphasizing that the system indeed follows the KZM at slow quenches, which is consistent with the behaviour of $\hat{t}$ and $\varepsilon(\hat{t})$ shown in Fig. 6(a) and 6(b). Notice however that as we approach $\tau_{q,c}$, marked by the vertical dashed lines, $N_{d}$ starts to fluctuate, with the fluctuation becoming more significant as $\varepsilon_{f}\rightarrow 0$. This behaviour is akin to the pre-saturation regime observed for closed systems under finite quench protocols [34]. As to whether this regime persists in the presence of quantum fluctuation remains an open question. We finally observe the saturation of $N_{d}$ as $\tau_{q}\rightarrow 0$, signifying the breakdown of the KZM for small values of $\tau_{q}$.
Figure 10: (a-b) Calculated $vz$ from the $N_{d}$ and threshold-based transition time as a function of (a) $\varepsilon_{f}$ for $\kappa=0.1\omega$ and (b) $\kappa$ for $\varepsilon_{f}=1.25$. The remaining parameter is $J=0.1\omega$. (c-d) Estimated critical exponents (c) $v$ and (d) $v$ as a function of $\varepsilon_{f}$. The dashed lines correspond to the critical exponents of the Ising universality class [62, 63]. The remaining parameters are set to $\kappa=J=0.1\omega$.
Since we have shown that the KZM quantities $\hat{t}_{\mathrm{th}}$, $\varepsilon(\hat{t}_{\mathrm{th}})$, and $N_{d}$ follows the predicted KZM scaling, for completeness, we now estimate the critical exponents of the open DLM from the power-law exponents of the of these quantities. We do this by assuming that $\hat{t}_{\mathrm{th}}$ and $N_{d}$ follows a generic power-law scaling,
@@
within the quench time interval $\omega\tau_{q}=\omega\tau_{q,c}$ and $\omega\tau_{q}=10^{4}$. From these equations, we can infer from Eq. (1) and Eq. (27) that $\alpha$ and $\beta$ are related to critical exponents $v$ and $z$ by the relations
@@
As for the behaviour of the threshold-based and defect-based $vz$ as a function of $\kappa$, we can see in Fig. 10(b) that both values of $vz$ decrease as $\kappa\rightarrow 0$. In particular, the $vz$ for both cases approaches the experimental value of $vz$ for the open Dicke model for $\kappa=1.0\omega$ [27]. This result implies that the system’s dissipation modifies the critical exponents of the system, which is consistent with the predictions in Refs. [21] and [22]. Without any specific analytical prediction on how $\kappa$ modifies the effective critical exponent of the system, we cannot assign a universality class for the open DLM that may apply to any arbitrary dissipation strength.

### v2

@@ V Summary and Discussion
In this paper, we extend the Kibble-Zurek mechanism to open systems under a finite quench and report an intermediate regime separating the breakdown and validity of the KZM at fast and slow quench timescales, respectively. This regime manifests as a continuation of the transition time’s KZM power-law scaling at $\tau_{q}$, where the system appears to relax after the quench has terminated. As we have shown using a coupled oscillator system, this phenomenon results from the system’s relaxation towards the global minimum of its potential due to dissipation. This mechanism effectively hides the system’s crossover to the adiabatic regime, only to be revealed once the system reaches the arbitrary threshold of the order parameter.
Using the open DM and the open DLM, we have also demonstrated that the delayed KZM is a generic feature of open systems under finite quenches that can be mapped onto a coupled oscillator system. Furthermore, we have shown that the signatures of the delayed KZM, specifically the size of the quench interval where the delayed KZM regime is observed, become more prominent for small values of $\varepsilon_{f}$, highlighting the dissipative and near-critical nature of this phenomenon. We have discussed the implications of the delayed KZM in the context of the threshold-based criterion typically used in experiments to measure the transition time and proposed an alternative method to measure $\hat{t}$. Our proposed method only relies on the spatialtemporal information of an appropriate order parameter, such as the defect number and phase information of the system’s bosonic modes, thus providing a more sensitive tool for detecting phase transitions.
Our results extend the notion of the KZM to dissipative systems with finite quench protocols beyond the limits of slow and rapid quenches. It also provides a framework on how the manifestation of the KZM can be altered in experimental protocols, wherein limitations in measuring the true AI crossover become more relevant. Since our results are all in the mean-field level, a natural extension of our paper is to verify whether the delayed KZM would survive in the presence of quantum fluctuations. It would also be interesting to test the signatures of the delayed KZM in the quantum regime of the open DM and the open DLM, and further explore their universality classes beyond the mean-field level. These extensions can be readily done in multiple platforms, including, but not limited to, cavity-QED setups [56, 57, 58, 28, 41], nitrogen-vacancy center ensembles [44, 59, 60, 61], cavity-magnon systems [46, 47], and photonic crystals [62].
Acknowledgements
@@ Appendix A Mean-field Equations of the Considered Systems
To obtain the mean-field equations of the systems considered in the main text, we consider the master equation for the expectation value of an arbitrary operator, $\hat{O}$,
@@
where $\hat{H}$ is the system’s Hamiltonian, and $\mathcal{D}\hat{O}=\sum_{\ell}\kappa_{\ell}\left(2\hat{L}_{\ell}^{\dagger}\hat{O}\hat{L}_{\ell}-\left\{\hat{L}^{\dagger}_{\ell}\hat{L}_{\ell},\hat{O}\right\}\right)$ is the dissipator, with $\hat{L}_{\ell}$ being the jump operators. We will also let $A=\left<\hat{A}\right>$ for notation convenience. Using this master equation, the mean-field equation of the open COS is
@@
As for the open Dicke model, its mean-field equations are
@@ Appendix B KZM exponents of the open Dicke Lattice model
where $D$ and $d$ are the dimensions of the system and the topological defects, respectively. To demonstrate that the open DLM satisfies the predicted KZM scaling for $N_{d}$, we present in Fig. 9 the number of phase slips present in the system for a given $\tau_{q}$ and $\varepsilon_{f}$. We can observe that for large $\tau_{q}$, $N_{d}$ follows a power-law scaling behavior with $\tau_{q}$, emphasizing that the system indeed follows the KZM at slow quenches, which is consistent with the behavior of $\hat{t}$ and $\varepsilon(\hat{t})$ shown in Figs. 6(a) and 6(b). Notice, however, that as we approach $\tau_{q,c}$, marked by the vertical dashed lines, $N_{d}$ starts to fluctuate, with the fluctuation becoming more significant as $\varepsilon_{f}\rightarrow 0$. This behavior is akin to the presaturation regime observed for closed systems under finite quench protocols [35]. As to whether this regime persists in the presence of quantum fluctuation remains an open question. We finally observe the saturation of $N_{d}$ as $\tau_{q}\rightarrow 0$, signifying the breakdown of the KZM for small values of $\tau_{q}$.
Since we have shown that the KZM quantities $\hat{t}_{\mathrm{th}}$, $\varepsilon(\hat{t}_{\mathrm{th}})$, and $N_{d}$ follow the predicted KZM scaling, for completeness, we now estimate the critical exponents of the open DLM from the power-law exponents of the of these quantities. We do this by assuming that $\hat{t}_{\mathrm{th}}$ and $N_{d}$ follows a generic power-law scaling,
@@
within the quench time interval $\omega\tau_{q}=\omega\tau_{q,c}$ and $\omega\tau_{q}=10^{4}$. From these equations, we can infer from Eqs. (1) and Eq. (27) that $\alpha$ and $\beta$ are related to critical exponents $v$ and $z$ by the relations
@@
Figure 10: (a), (b) Calculated $vz$ from the $N_{d}$ and threshold-based transition time as a function of (a) $\varepsilon_{f}$ for $\kappa=0.1\omega$ and (b) $\kappa$ for $\varepsilon_{f}=1.25$. The remaining parameter is $J=0.1\omega$. (c), (d) Estimated critical exponents (c) $v$ and (d) $z$ as a function of $\varepsilon_{f}$. The dashed lines correspond to the critical exponents of the Ising universality class [63, 64]. The remaining parameters are set to $\kappa=J=0.1\omega$.
@@
As for the behavior of the threshold-based and defect-based $vz$ as a function of $\kappa$, we can see in Fig. 10(b) that both values of $vz$ decrease as $\kappa\rightarrow 0$. In particular, the $vz$ for both cases approaches the experimental value of $vz$ for the open Dicke model for $\kappa=1.0\omega$ [28]. This result implies that the system’s dissipation modifies the critical exponents of the system, which is consistent with the predictions in Refs. [22] and [23]. Without any specific analytical prediction on how $\kappa$ modifies the effective critical exponent of the system, we cannot assign a universality class for the open DLM that may apply to any arbitrary dissipation strength.